How much 10.0 M HNO3 must be added to 1.00 L of a buffer that is 0.0100 M acetic acid and 0.100 M sodium acetate to reduce the pH to 5.35 ?he pKa of acetic acid is 4.75.
Answer #1

moles of acetic acid = 0.0100 mol

moles of sodium acetate = 0.100 mol

pH = 5.35

pKa = 4.75

pH = pKa + log [salt - C / acid + C]

5.35 = 4.75 + log [0.100 - C / 0.0100 + C]

[0.100 - C / 0.0100 + C] = 3.981

0.100 - C = 0.03981 +3.981 C

C = 0.01208

moles of HNO3 = 0.01208

10 x V = 0.01208

V = 0.001208 L

volume of HNO3 = 1.21 mL